# Accelerating Network Slice Placement with Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Cellular networks are increasingly realized through software-based entities, with core functions deployed as Virtual Network Functions (VNFs) on Commercial-off-the-Shelf (COTS) hardware. Network slicing has emerged as a key enabler of 5G by providing logically isolated Quality of Service (QoS) guarantees for diverse applications. With the adoption of cloud-native infrastructures, the placement of network slices across heterogeneous multi-cloud environments poses new challenges due to variable resource capabilities and slice-specific requirements. This paper introduces a modular framework for autonomous and near-optimal VNF placement based on a disaggregated Multi-Agent Reinforcement Learning (MARL) approach. The framework incorporates real traffic profiles to estimate slice resource demands and employs a MARL-based scheduler to minimize deployment cost while meeting QoS constraints. Experimental evaluation on a multi-cloud testbed shows a 19× speed-up compared to combinatorial optimization, with deployment costs within 7.8% of the optimal. While the method incurs up to 2.42× more QoS violations under high load, the trade-off provides significantly faster decision-making and reduced computational complexity. These results suggest that MARL-based approaches offer a scalable and cost-efficient solution for real-time network slice placement in heterogeneous infrastructures.
@@ I Introduction
Next-generation cellular networks are evolving to support various use cases, each with distinct Quality of Service (QoS) requirements. To accommodate this evolving diversity, “network slicing” has emerged as a key enabler, facilitating the creation of virtually isolated network segments over a shared physical infrastructure to support different use cases like URLLC (Ultra-Reliable Low-Latency Communication), eMBB (Enhanced Mobile Broadband), and mMTC (massive Machine Type Communications) [1, 2].
@@
The advent of network slicing is facilitated by the adoption of foundational technologies such as Network Function Virtualization (NFV) and Software Defined Networking (SDN) [3, 4]. Leveraging NFV and SDN, the proprietary Physical Network Functions (PNFs) that made up the legacy 4G Radio Access Network (RAN) and Core Network (CN), are now deployed as Virtual Network Functions (VNFs) on Commercial-off-the-Shelf (COTS) hardware in the form of finer-grained microservices. A network slice is ultimately formed by service chaining a set of VNFs to create an isolated network segment that extends from the RAN to the CN [2].
@@
To capitalize on this software-driven approach, network operators are partnering with cloud providers to advance cloud-based 5G deployments [5, 6]. This shift enables the dynamic instantiation and migration of network slice VNFs across a geographically distributed cloud infrastructure. By leveraging this flexibility, operators can tailor deployment strategies to accommodate the diverse Service Level Agreements (SLAs) of different use cases. Specifically, operators can strategically position latency-sensitive VNFs closer to the RAN at edge locations [7, 8], despite higher computational costs, while hosting less critical functions more cost-effectively in regional or central clouds [2, 9, 10]. Modern multi-cloud infrastructures provide a wide range of computing, networking, and storage capabilities. With infrastructure heterogeneity, coupled with diverse service requirements in terms of latency, bandwidth, and reliability, strategic VNF placement decisions become essential for optimizing resource use and supporting varied application demands in next-generation networks.
@@
Motivation. The problem of network slice placement is widely recognized in the literature as a variant of the Virtual Network Embedding (VNE) problem, which is proven to be NP-hard  [9]. Prior work has explored Integer Linear Programming (ILP)-based methods to achieve optimality [11, 12, 13], but these approaches do not scale and are computationally intensive, making them impractical for real-time decisions. Heuristic-based strategies have been proposed to improve efficiency, yet they often fail to achieve solutions close to the optimum, as they tend to get stuck in local optima  [14, 15]. More recently, Reinforcement Learning (RL) has gained traction as a promising approach for optimal slice placement through trial and error procedures [16, 17]. However, existing RL solutions predominantly rely on monolithic agent implementations that struggle to capture resource and service constraints of different slice types [18, 19, 20], lack real-world validation [20, 21], and scale poorly in large deployments [19, 20].
@@
  • We analyze the resource demands of key user plane VNFs by replaying the captured traffic, enabling accurate resource estimation and slice placement optimization through enhancements to the open-source OpenAirInterface (OAI) 5G software stack [22].
@@
  • To foster reproducibility and enable further research, we plan to publicly release all artifacts, including the collected datasets, source code, and deployment scripts.
@@ II BACKGROUND
Radio Access Network: Unlike previous generations, the 5G RAN embraces a disaggregated virtualized architecture that enhances efficiency and scalability by allowing network operators to dynamically allocate resources based on service demands [23]. This is achieved by virtualizing baseband processing functions, enabling flexible deployment on cloud infrastructures. The RAN is composed of three main units: the Radio Unit (RU), which handles radio transmission, reception, and antenna processing; the Distributed Unit (DU), responsible for lower-layer processing such as error correction and resource allocation; and the Centralized Unit (CU), which manages higher-layer functions [23].
Core Network: The 5G CN is designed with a service-based, cloud-native architecture to ensure modularity, scalability, and support for network slicing [24]. It is composed of several control plane VNFs that operate on a consumer-producer model, communicating with each other via standardized REST APIs. At the heart of the 5G Core is the Access and Mobility Management Function (AMF), which acts as the communication hub between the RAN, UE, and other VNFs, providing essential signaling and mobility management services. The Authentication Server Function (AUSF) provides authentication credentials to the AMF, working in conjunction with the Unified Data Management (UDM) and Unified Data Repository (UDR) to implement the 5G Authentication and Key Agreement (AKA) procedure. Within the CN, the Session Management Function (SMF) and User Plane Function (UPF) serve as the control and user plane anchors, respectively. The UPF tunnels user traffic toward an external DN node, which represents real-world application servers. Finally, the Network Repository Function (NRF) facilitates initial discovery and communication among VNFs.
@@
Network Slicing. A network slice is an end-to-end logical network instance formed by service chaining a set of VNFs, including both control and user plane components, to meet specific application requirements. Each VNF has distinct resource requirements, and is deployed across a virtualized infrastructure. While network slicing enables unprecedented flexibility by allowing operators to create isolated, end-to-end virtual networks, it also introduces complex challenges in orchestration, resource allocation, and slice placement optimization [2]. Several standardization efforts have been proposed to define and manage the lifecycle of network slices. Among them, the ETSI NFV architecture provides a comprehensive framework for deploying VNFs, and automating slice operations [25]. At its core, the Network Function Virtualization Orchestrator (NFVO) coordinates slice instantiation, scaling, and lifecycle management. The Virtualized Infrastructure Manager (VIM) oversees computing, storage, and networking resources, translating NFVO decisions into infrastructure-level actions. Beneath these layers, the Network Functions Virtualization Infrastructure (NFVI) unifies physical and virtual resources across edge, distributed, and central cloud domains, providing the foundation for hosting network slices.
@@ III-A User Plane Traffic Capturing Block
The UPTCB is the first building block of the SlicePilot framework, designed to collect, preprocess, and profile over-the-air user plane traffic, enabling realistic slice performance analysis and optimization. Unlike synthetic traffic generators such as iPerf3 [26], which offer controlled traffic generation but lack configurability for complex application-specific patterns, UPTCB builds a structured repository of preprocessed flows aligned with the three primary 5G slice types: eMBB for bandwidth-intensive applications, URLLC for ultra-low latency services, and mMTC for massive IoT deployments. By replaying these traffic profiles through the user plane path, as illustrated in Figure 2, UPTCB enables accurate VNF resource estimation under real-world conditions, supporting informed slice placement and scaling. For interested readers, details on the traffic collection methodology are provided in Section §IV.
@@ III-C Slice Placement Block
The final block of the SlicePilot framework is a fully customized MARL scheduler, designed to optimize slice deployment within a heterogeneous multi-cloud infrastructure. It employs a disaggregated agent design to accelerate convergence, improve stability, and simplify decision-making by assigning each agent to manage slice requests of the same type. This specialization reduces the state-action space each agent needs to explore, making slice placement more efficient and scalable compared to monolithic single-agent approaches [17, 27, 28]. Operating within the NFVO, the scheduler computes placement decisions and relays them to the VIM for slice deployment and instantiation within the NFVI. The overall architecture of the SPB is illustrated in Figure 6.
@@
Slice-Wise Problem Decomposition. We decompose the optimization problem described in (6) into a set of smaller optimization problems, each corresponding to a specific slice type $z\in Z$. These sub-problems can be solved independently based on slice request arrival times, i.e., the first slice request is processed first, followed by subsequent requests in their respective order of arrival. We model each sub-problem as a Markov Decision Process (MDP) [29], viewing it as a sequential decision-making process where VNFs are placed one at a time until all required VNFs are deployed. Each MDP state $s_{t}\in S$ is described as follows:
@@
A key challenge in this formulation is how agents retain and utilize past experiences. Classical RL methods often rely on tabular approaches to store and update experiences, but these become impractical in large state-action spaces due to limited generalization [30]. In our case, the state-action space per agent grows exponentially as $\mathcal{O}(|V|\cdot|M|\cdot|Q_{z}|\cdot|M|^{|V|})$, necessitating scalable learning techniques for efficient generalization. Function approximation has been explored to address this issue [31], with Deep Neural Networks (DNNs) emerging as a promising solution due to their success in recent applications [27, 28]. Motivated by this, we employ DNNs as value function approximators within our agents to cope with the high-dimensional state-action space.
Deep Q-Learning for Value Function Approximation. To enable efficient decision-making in each agent, we adopt Deep Q-Learning (DQL), an RL approach that integrates Q-learning with DNNs to generalize over high-dimensional state-action spaces [32, 33]. The goal is to learn a Q-function that maps each state-action pair $(s,a)$ to the expected cumulative reward obtained by taking action $a$ in state $s$ and following a policy thereafter, formally defined as $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$.
@@
where $s^{\prime}$ is the next state, $a^{\prime}$ is the next action, and the expectation is taken over the transition distribution $P(s^{\prime}|s,a)$. To approximate $Q^{*}(s,a)$, we use a DNN parameterized by $\theta$, denoted as $Q_{\theta}(s,a)$. The agent iteratively updates this network through a value iteration process that refines its reward estimates over time [32]. This is achieved by minimizing the difference between the predicted and target Q-values using a mean squared error loss:
@@
    Target Network. We maintain a delayed version of the Q-function, $Q_{\theta^{\prime}}(s,a)$ [33], to compute the target value more reliably. The target in (15) is modified as:
@@
    Double Q-Learning. To mitigate the overestimation bias caused by using the same Q-function for both action selection and evaluation, we use Double Q-learning [34]. The action that maximizes the Q-value is selected using the current network:
@@ IV EXPERIMENTATION ENVIRONMENT
User Plane Traffic Collection. For the traffic collection campaign, we leveraged our in-lab over-the-air 5G standalone testbed, which operated in the n78 TDD band with a 40 MHz bandwidth and utilized a USRP B210 SDR as the RF front end, as illustrated in Figure 7. The baseband processing and core network VNFs were deployed as Docker containers using the latest OAI release [22]. A Google Pixel 7 mobile device served as the primary UE for most network scenarios. Additionally, to connect external devices, we tethered the Pixel to an OpenWRT router acting as a gateway to our 5G network. We considered two external devices: an AR/VR headset (Microsoft HoloLens 2) and a Govee WiFi Thermometer Hygrometer H5179 smart humidity and temperature IoT sensor.
Building on this setup, we conducted a series of traffic collection experiments targeting three slice types: URLLC, eMBB, and mMTC. For each application scenario, we launched the base station (gNB), verified its connection to the core network, and connected the UE by disabling airplane mode. Once the SMF assigned an IP address, we captured traffic on the gNB host using tcpdump [35], filtering by the UE’s IP to isolate user-plane packets and exclude control signaling. Each scenario was recorded for 4 hours to ensure sufficient traffic volume and diversity. The application use cases were grouped by slice type: URLLC included VoIP (Messenger, WhatsApp), online gaming (Fortnite, PUBG), web browsing, and AR/VR (Microsoft Dynamics 365 Remote Assist); eMBB included media streaming and downloads (YouTube, TikTok), social media (Facebook, Instagram), and video conferencing (Messenger, Skype, Zoom); mMTC focused on IoT traffic.
@@ V-A Configurations and Settings
Infrastructure Configurations. We consider $M=3$ heterogeneous cloud infrastructures—edge, distributed, and central—and set their resource capacities as follows: $C_{0}=$ (16 CPU cores, 16 GiB RAM), $C_{1}=$ (32 CPU cores, 32 GiB RAM), and $C_{2}=$ (64 CPU cores, 64 GiB RAM), respectively. Each infrastructure hosts virtual machines instantiated using predefined flavors: (1 CPU core, 1 GiB RAM), (2 CPU cores, 2 GiB RAM), and (4 CPU cores, 4 GiB RAM). We select cost values to reflect the hourly charge for on-demand reservation of 1 CPU core and 1 GiB RAM, set as $\lambda_{0}=0.010$ \$/h at the edge, $\lambda_{1}=0.005$ \$/h at the distributed cloud, and $\lambda_{2}=0.001$ \$/h at the central cloud, following commercial pricing trends [36]. Inter-cloud latencies are modeled as normal distributions, following values reported in [37]: edge-to-distributed latency $D_{0,1}^{\text{link}}\sim\mathcal{N}(0.5,0.1^{2})\,\text{ms}$, and distributed-to-central $D_{1,2}^{\text{link}}\sim\mathcal{N}(20,1^{2})\,\text{ms}$. Finally, latency to external networks depends on UPF placement—5 ms at the edge, 7.5 ms in the distributed cloud, and 10 ms in the central cloud.

### v2

# SlicePilot: Demystifying Network Slice Placement in Heterogeneous Cloud Infrastructures

@@ Abstract
Cellular networks are comprised of software-based entities, with main functions encapsulated as Virtual Network Functions (VNFs) deployed on Commercial-off-the-Shelf (COTS) hardware. As a key enabler of 5G, network slicing offers logically isolated Quality of Service (QoS) for diverse use cases. With the transition to cloud-native infrastructures, optimizing network slice placement across multi-cloud environments remains challenging due to heterogeneous resource capabilities and varying slice-specific demands. This paper presents SlicePilot, a modular framework that enables autonomous and near-optimal VNF placement using a disaggregated Multi-Agent Reinforcement Learning (MARL) approach. SlicePilot collects real-world traffic profiles to estimate resource needs for each slice type. These estimates guide a MARL-based scheduler that minimizes deployment costs while satisfying QoS constraints. We evaluate SlicePilot on a multi-cloud testbed and demonstrate a 19× speed-up over combinatorial optimization methods, while keeping deployment costs within 7.8% of the optimal. Although SlicePilot results in 2.42× more QoS violations under high-load conditions, this trade-off is offset by faster decision-making and reduced computational overhead. Overall, SlicePilot delivers a scalable, cost-efficient solution for network slice placement, making it suitable for real-time deployments where responsiveness and efficiency are critical.
@@ I Introduction
Next-generation cellular networks are evolving to support various use cases, each with distinct Quality of Service (QoS) requirements. To accommodate this evolving diversity, “network slicing” has emerged as a key enabler, facilitating the creation of virtually isolated network segments over a shared physical infrastructure to support different use cases like URLLC (Ultra-Reliable Low-Latency Communication), eMBB (Enhanced Mobile Broadband), and mMTC (massive Machine Type Communications) [surv1, surv4].
@@
The advent of network slicing is facilitated by the adoption of foundational technologies such as Network Function Virtualization (NFV) and Software Defined Networking (SDN) [etsi_nfv_whitepaper, onf_sdn_whitepaper]. Leveraging NFV and SDN, the proprietary Physical Network Functions (PNFs) that made up the legacy 4G Radio Access Network (RAN) and Core Network (CN), are now deployed as Virtual Network Functions (VNFs) on Commercial-off-the-Shelf (COTS) hardware in the form of finer-grained microservices. A network slice is ultimately formed by service chaining a set of VNFs to create an isolated network segment that extends from the RAN to the CN [surv4].
@@
To capitalize on this software-driven approach, network operators are partnering with cloud providers to advance cloud-based 5G deployments [5GCore5G67online, 5GCoreon75online]. This shift enables the dynamic instantiation and migration of network slice VNFs across a geographically distributed cloud infrastructure. By leveraging this flexibility, operators can tailor deployment strategies to accommodate the diverse Service Level Agreements (SLAs) of different use cases. Specifically, operators can strategically position latency-sensitive VNFs closer to the RAN at edge locations [5GEdgeCo6online, LowLaten55online], despite higher computational costs, while hosting less critical functions more cost-effectively in regional or central clouds [surv4, surv6, surv7]. Modern multi-cloud infrastructures provide a wide range of computing, networking, and storage capabilities. With infrastructure heterogeneity, coupled with diverse service requirements in terms of latency, bandwidth, and reliability, strategic VNF placement decisions become essential for optimizing resource use and supporting varied application demands in next-generation networks.
@@
Motivation. The problem of network slice placement is widely recognized in the literature as a variant of the Virtual Network Embedding (VNE) problem, which is proven to be NP-hard  [surv6]. Prior work has explored Integer Linear Programming (ILP)-based methods to achieve optimality [opt1, opt2, opt3], but these approaches do not scale and are computationally intensive, making them impractical for real-time decisions. Heuristic-based strategies have been proposed to improve efficiency, yet they often fail to achieve solutions close to the optimum, as they tend to get stuck in local optima  [opt5, opt4]. More recently, Reinforcement Learning (RL) has gained traction as a promising approach for optimal slice placement through trial and error procedures [rl4, rl6]. However, existing RL solutions predominantly rely on monolithic agent implementations that struggle to capture resource and service constraints of different slice types [rl1, rl2, rl3], lack real-world validation [rl3, testbed2], and scale poorly in large deployments [rl2, rl3].
@@
  • We analyze the resource demands of key user plane VNFs by replaying the captured traffic, enabling accurate resource estimation and slice placement optimization through enhancements to the open-source OpenAirInterface (OAI) 5G software stack [oai].
@@ II BACKGROUND
Radio Access Network: Unlike previous generations, the 5G RAN embraces a disaggregated virtualized architecture that enhances efficiency and scalability by allowing network operators to dynamically allocate resources based on service demands [polese2023understanding]. This is achieved by virtualizing baseband processing functions, enabling flexible deployment on cloud infrastructures. The RAN is composed of three main units: the Radio Unit (RU), which handles radio transmission, reception, and antenna processing; the Distributed Unit (DU), responsible for lower-layer processing such as error correction and resource allocation; and the Centralized Unit (CU), which manages higher-layer functions [polese2023understanding].
Core Network: The 5G CN is designed with a service-based, cloud-native architecture to ensure modularity, scalability, and support for network slicing [3GPP_TS_23_501]. It is composed of several control plane VNFs that operate on a consumer-producer model, communicating with each other via standardized REST APIs. At the heart of the 5G Core is the Access and Mobility Management Function (AMF), which acts as the communication hub between the RAN, UE, and other VNFs, providing essential signaling and mobility management services. The Authentication Server Function (AUSF) provides authentication credentials to the AMF, working in conjunction with the Unified Data Management (UDM) and Unified Data Repository (UDR) to implement the 5G Authentication and Key Agreement (AKA) procedure. Within the CN, the Session Management Function (SMF) and User Plane Function (UPF) serve as the control and user plane anchors, respectively. The UPF tunnels user traffic toward an external DN node, which represents real-world application servers. Finally, the Network Repository Function (NRF) facilitates initial discovery and communication among VNFs.
@@
Network Slicing. A network slice is an end-to-end logical network instance formed by service chaining a set of VNFs, including both control and user plane components, to meet specific application requirements. Each VNF has distinct resource requirements, and is deployed across a virtualized infrastructure. While network slicing enables unprecedented flexibility by allowing operators to create isolated, end-to-end virtual networks, it also introduces complex challenges in orchestration, resource allocation, and slice placement optimization [surv4]. Several standardization efforts have been proposed to define and manage the lifecycle of network slices. Among them, the ETSI NFV architecture provides a comprehensive framework for deploying VNFs, and automating slice operations [etsi_nfveve012v311]. At its core, the Network Function Virtualization Orchestrator (NFVO) coordinates slice instantiation, scaling, and lifecycle management. The Virtualized Infrastructure Manager (VIM) oversees computing, storage, and networking resources, translating NFVO decisions into infrastructure-level actions. Beneath these layers, the Network Functions Virtualization Infrastructure (NFVI) unifies physical and virtual resources across edge, distributed, and central cloud domains, providing the foundation for hosting network slices.
@@ III-A User Plane Traffic Capturing Block
The UPTCB is the first building block of the SlicePilot framework, designed to collect, preprocess, and profile over-the-air user plane traffic, enabling realistic slice performance analysis and optimization. Unlike synthetic traffic generators such as iPerf3 [iperf3], which offer controlled traffic generation but lack configurability for complex application-specific patterns, UPTCB builds a structured repository of preprocessed flows aligned with the three primary 5G slice types: eMBB for bandwidth-intensive applications, URLLC for ultra-low latency services, and mMTC for massive IoT deployments. By replaying these traffic profiles through the user plane path, as illustrated in Figure 2, UPTCB enables accurate VNF resource estimation under real-world conditions, supporting informed slice placement and scaling. For interested readers, details on the traffic collection methodology are provided in Section §IV.
@@ III-C Slice Placement Block
The final block of the SlicePilot framework is a fully customized MARL scheduler, designed to optimize slice deployment within a heterogeneous multi-cloud infrastructure. It employs a disaggregated agent design to accelerate convergence, improve stability, and simplify decision-making by assigning each agent to manage slice requests of the same type. This specialization reduces the state-action space each agent needs to explore, making slice placement more efficient and scalable compared to monolithic single-agent approaches [rl6, rl7, rl8]. Operating within the NFVO, the scheduler computes placement decisions and relays them to the VIM for slice deployment and instantiation within the NFVI. The overall architecture of the SPB is illustrated in Figure 6.
@@
Slice-Wise Problem Decomposition. We decompose the optimization problem described in (6) into a set of smaller optimization problems, each corresponding to a specific slice type $z\in Z$. These sub-problems can be solved independently based on slice request arrival times, i.e., the first slice request is processed first, followed by subsequent requests in their respective order of arrival. We model each sub-problem as a Markov Decision Process (MDP) [mdp], viewing it as a sequential decision-making process where VNFs are placed one at a time until all required VNFs are deployed. Each MDP state $s_{t}\in S$ is described as follows:
@@
A key challenge in this formulation is how agents retain and utilize past experiences. Classical RL methods often rely on tabular approaches to store and update experiences, but these become impractical in large state-action spaces due to limited generalization [sutton2018reinforcement]. In our case, the state-action space per agent grows exponentially as $\mathcal{O}(|V|\cdot|M|\cdot|Q_{z}|\cdot|M|^{|V|})$, necessitating scalable learning techniques for efficient generalization. Function approximation has been explored to address this issue [bertsekas], with Deep Neural Networks (DNNs) emerging as a promising solution due to their success in recent applications [rl7, rl8]. Motivated by this, we employ DNNs as value function approximators within our agents to cope with the high-dimensional state-action space.
Deep Q-Learning for Value Function Approximation. To enable efficient decision-making in each agent, we adopt Deep Q-Learning (DQL), an RL approach that integrates Q-learning with DNNs to generalize over high-dimensional state-action spaces [qlearning, dqn]. The goal is to learn a Q-function that maps each state-action pair $(s,a)$ to the expected cumulative reward obtained by taking action $a$ in state $s$ and following a policy thereafter, formally defined as $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$.
@@
where $s^{\prime}$ is the next state, $a^{\prime}$ is the next action, and the expectation is taken over the transition distribution $P(s^{\prime}|s,a)$. To approximate $Q^{*}(s,a)$, we use a DNN parameterized by $\theta$, denoted as $Q_{\theta}(s,a)$. The agent iteratively updates this network through a value iteration process that refines its reward estimates over time [qlearning]. This is achieved by minimizing the difference between the predicted and target Q-values using a mean squared error loss:
@@
    Target Network. We maintain a delayed version of the Q-function, $Q_{\theta^{\prime}}(s,a)$ [dqn], to compute the target value more reliably. The target in (15) is modified as:
@@
    Double Q-Learning. To mitigate the overestimation bias caused by using the same Q-function for both action selection and evaluation, we use Double Q-learning [ddqn]. The action that maximizes the Q-value is selected using the current network:
@@ IV EXPERIMENTATION ENVIRONMENT
User Plane Traffic Collection. For the traffic collection campaign, we leveraged our in-lab over-the-air 5G standalone testbed, which operated in the n78 TDD band with a 40 MHz bandwidth and utilized a USRP B210 SDR as the RF front end, as illustrated in Figure 7. The baseband processing and core network VNFs were deployed as Docker containers using the latest OAI release [oai]. A Google Pixel 7 mobile device served as the primary UE for most network scenarios. Additionally, to connect external devices, we tethered the Pixel to an OpenWRT router acting as a gateway to our 5G network. We considered two external devices: an AR/VR headset (Microsoft HoloLens 2) and a Govee WiFi Thermometer Hygrometer H5179 smart humidity and temperature IoT sensor.
Building on this setup, we conducted a series of traffic collection experiments targeting three slice types: URLLC, eMBB, and mMTC. For each application scenario, we launched the base station (gNB), verified its connection to the core network, and connected the UE by disabling airplane mode. Once the SMF assigned an IP address, we captured traffic on the gNB host using tcpdump [tcpdump], filtering by the UE’s IP to isolate user-plane packets and exclude control signaling. Each scenario was recorded for 4 hours to ensure sufficient traffic volume and diversity. The application use cases were grouped by slice type: URLLC included VoIP (Messenger, WhatsApp), online gaming (Fortnite, PUBG), web browsing, and AR/VR (Microsoft Dynamics 365 Remote Assist); eMBB included media streaming and downloads (YouTube, TikTok), social media (Facebook, Instagram), and video conferencing (Messenger, Skype, Zoom); mMTC focused on IoT traffic.
@@ V-A Configurations and Settings
Infrastructure Configurations. We consider $M=3$ heterogeneous cloud infrastructures—edge, distributed, and central—and set their resource capacities as follows: $C_{0}=$ (16 CPU cores, 16 GiB RAM), $C_{1}=$ (32 CPU cores, 32 GiB RAM), and $C_{2}=$ (64 CPU cores, 64 GiB RAM), respectively. Each infrastructure hosts virtual machines instantiated using predefined flavors: (1 CPU core, 1 GiB RAM), (2 CPU cores, 2 GiB RAM), and (4 CPU cores, 4 GiB RAM). We select cost values to reflect the hourly charge for on-demand reservation of 1 CPU core and 1 GiB RAM, set as $\lambda_{0}=0.010$ \$/h at the edge, $\lambda_{1}=0.005$ \$/h at the distributed cloud, and $\lambda_{2}=0.001$ \$/h at the central cloud, following commercial pricing trends [aws_ec2_pricing]. Inter-cloud latencies are modeled as normal distributions, following values reported in [atalay2023first]: edge-to-distributed latency $D_{0,1}^{\text{link}}\sim\mathcal{N}(0.5,0.1^{2})\,\text{ms}$, and distributed-to-central $D_{1,2}^{\text{link}}\sim\mathcal{N}(20,1^{2})\,\text{ms}$. Finally, latency to external networks depends on UPF placement—5 ms at the edge, 7.5 ms in the distributed cloud, and 10 ms in the central cloud.